# On the foundations and applications of Lorentz-Finsler Geometry

### v1

@@ 5. Real world applications
    Finsler spacetime techniques can solve technical issues about propagation which are much more involved from a purely Finslerian viewpoint (for example, see the mention to [92, Lemma 4.7] in §5.2.1 below).
@@ 5.1.1. An increasingly general Fermat principle
Figure 11. The first and last causal curves from $p$ to $\alpha$ (in green) must be lightlike geodesics by causality considerations. They are also some of the possible critical points of the arrival functional.
@@
In the relativistic setting, the problem was posed by Kovner [108] and solved rigurously by Perlick [141]. This author also noticed that the existence of the global time $t$ was irrelevant and, indeed the parameter of $\alpha$ for any regular parametrization (i.e., smooth with non-vanishing velocity) could play its role for the functional $\mathcal{T}$. Moreover, as the computations of critical points only uses variations of the arriving curve, global hypotheses in $\mathcal{C}$ are not required. Perlick also made a first extension of Fermat principle to the Lorentz-Finsler case [143].
Perlick and Piccione [144] also extended the result to the case when the source of the lightlike curves is a spacelike submanifold $P$ (instead of a point $p$) and the curves arrive at a timelike submanifold (now actually endowed with a time function in order to redefine properly $\mathcal{T}$). In the setting of SSTK spacetimes above, Caponio et al. [39] considered also the case when the arrival curve $\alpha$ is not timelike, noticing that just the non-orthogonality between $\alpha$ and the incident curve $\gamma$ is sufficient to establish Fermat principle (notice that this condition is automatically satisfied when $\alpha$ is timelike), see [39, Th. 7.4].
@@ 5.1.2. Zermelo navigation
If $\mathcal{C}$ is a SSTK structure [^20] The following assertions can be extended to Finsler SSTK spacetimes, taking into account footnote 16. then its whole causality is encoded in a wind Finsler structure $\Sigma$ on a slice $S=\{t=0\}$ through $p$ of its time $t$. Thus, the general conditions which ensure causal simplicity or global hyperbolicity for $\mathcal{C}$ can be expressed accurately in terms of $\Sigma$ as explained in §4.1. If $\alpha$ is an integral curve of the Killing vector field $K$ intersecting $S$ at a point $q$, then the existence of Zermelo solutions is equivalent to the existence of minimizing geodesics $\rho$ for the wind Finsler structure from $p$ to $q$. If, additionally, $\alpha$ is parametrized by the time $t$, then the arrival times is the length of $\rho$. In case that the SSTK is a Finsler standard stationary spacetime, the wind Finsler structure $\Sigma$ becomes just a Finsler metric and $\rho$ a usual geodesic. Moreover, as the integral curves of $K$ are timelike, the Fermat principle holds with no restriction about orthogonality at the arrival point in $\alpha$. From the geometric viewpoint, this generalizes widely the previous study by Perlick about the stationary case in [142], where the possibility of a reduction in one dimension less for Fermat principle considered.
@@ 5.1.3. Snell law
The basic Snell law for (anisotropic) cones is ([95, Th. 4.3], see Fig. 14):
@@
where $u$ (resp $v$) is the incident (resp. refracted) $\mathcal{C}_{1}$-lightlike (resp. $\mathcal{C}_{2}$-lightlike) direction at $o\in\eta$, $H_{u}^{1}$ (resp. $H_{v}^{2}$) is the hyperplane tangent to $\mathcal{C}_{1}$ at $u$ (resp. to $\mathcal{C}_{2}$ at $v$) and Snell’s formula states that the intersections $\Pi_{u}^{1},\Pi_{v}^{2}$ of these hyperplanes with the tangent space to the interface $T_{0}\eta$ must be equal. There is, however, a number of issues in relation to (22) and Fig. 14:
@@
  • The orientations in these hyperplanes selected by the side where the corresponding cone lives must be consistent, as in Fig. 14. This consistency is the additional hypothesis to ensure local minimization for Fermat principle and, then, that we obtain a Snell geodesic i.e. a locally horismotic curve.
    Notice that, if the blue cone lived on the other side of $H_{v}^{2}$, then the broken lightlike geodesic could not minimize the arrival time (and could not be used to compute the causal future of $p$).
  • Classical effects as reflection and total reflection also appear as a consequence of Fermat problem.
@@
    For example, when $\eta$ is spacelike for $Q_{1}$ and $Q_{2}$ the reflection is forbidden. Moreover, if the cones point out in consistent oriented directions at $\eta$, two refracted trajectories would appear, see Fig. 15. In this case, only one of them is minimizing (a fact determined again by orientations) providing then a Snell geodesic. Otherwise, no refracted one can appear, see [^21] In these figures, past cones (which were not considered in $\mathcal{C}$) are depicted and filled for an easier comparison with usual relativistic diagrams. Fig. 16.
@@
All these possibilities are analyzed in [95, §5], as well as the corresponding causality in such discontinuous media [95, §6]. About the latter, it is worth pointing out the higher technical difficulty of the Lorentz-Finsler case in comparison with the Lorentz one, stressed in the proof of [95, Lemma 6.6].
@@ 5.2.1. General wave propagation and wildfires
Reasoning as for Zermelo navigation in §5.1.2, let $(M,\mathcal{C})$ be a cone structure modelling the velocity of wave propagation in each event and direction permitted by a medium. Notice that, in the case of Cosmology, the medium is vacuum, but in the case of a problem in classical Mechanics, as the propagation of a wildfire or a sound wave, a cone triple for $(\Omega,T,F)$ on $\mathcal{C}$ appears naturally (namely, $T$ models the observers “at rest” on Earth, $\Omega$ is $dt$, with $t$ the Newtonian time, and $F$ is then determined by $\mathcal{C}$).
Focusing on a wildfire, it will start at a bounded region $R$ which can be modelled as a compact open subset with boundary $S=\partial R$ of the (spacelike) hypersurface $t=0$. Then, the evolved burned region is equal to its $\mathcal{C}$-causal future $J^{+}(R)$. The spacetime frontfire will be its boundary in $t>0$, which matches the horismos $E^{+}(S)$. The frontfire at each instant $t_{0}>0$ is $E^{+}(S)\cap\{t=t_{0}\}$ (see Fig. 17). Now, Lorentz-Finsler geometry applies to show that $E^{+}(S)$ is composed by the cone geodesics starting at $S$ with initial velocity pointing outside $\mathds{R}$. Notice that at each point $p\in S$ there are exactly two lightike directions orthogonal to $S$, the second one pointing inwards $R$.
This general viewpoint of anisotropic wave propagation is developed in [92]. It includes the short time existence of solutions [92, Theorem 4.8]. This result is more involved now than in the purely Lorentz case (see [92, Lemma 4.7]), however, it had not been carried out previously (i.e., with a non-spacetime approach) as far as we know.
The particular case of wildfires is analyzed detailedly in [94]. This includes a comparison with actual wildfire monitorization, which is based in rough approximations to the wave equation. As emphasized therein [94, §5.2], focal and cut points along cone geodesics become points where wildfires are more intense and firefighters can get trapped.
@@ 5.2.2. From seisms to Numerics and Particle Physics
The modelization of seisms is similar to wildfires. However, it is natural to consider discontinuous layers for seisms and, thus, a systematic use of Snell formula. Notice that the interface between different layers yields a timelike interface $\eta$ where the basic Snell formula (22) is enough. Indeed, typically, the interface will be “constant in time” so that its spacetime evolution is trivial [^22] Certainly, the effect of the seism itself could modify the interface, however, this would be negligible for the computation of the sismic wavefront. Moreover, in any case, it is not expected that the velocity of the movement of the interface could be compared with the velocity of seism propagation so that (22) would remain..
Anyway, the use of Snell formula can go far beyond the scope of seims, and required other cases explained below (22):
@@
    Computer aided modelizations of these classical systems must use discretizations. In a natural way, this modelization yield a grid where each cell would contains a cone (Fig. 18). Then, the propagation of the wave between each two grids will be controlled by the basic Snell formula (22).
@@
    Such a discretization may be useful for Numerical Relativity. The usual way to proceed in this field is considering at each cell the Lorentzian metric at one of its points. Notice, however, that such a Lorentzian metric is determined by the cone structure $\mathcal{C}$ up to a conformal factor. Thus, the separate study of $\mathcal{C}$ as above and this factor might be more efficient.
    Notice that, for this purpose, the interface would be generically timelike or spacelike. The first case is essentially governed by Snell formula (22). For spacelike interfaces, the possibility of double refraction (as in 15) appears naturally. Then, the possibility to distinguish the Snell geodesics by using orientations would be useful.
    Nevertheless, the the case of no refraction depicted Fig. 16 would mean either that the grid approximation is overly rough or that a singularity appears.
@@ 6.1.1. Very Special and General Relativity
where $\eta_{\mu\nu}$ is the Minkowski metric, $n^{\mu}$ is a fixed lightlike direction ($\eta_{\mu\nu}n^{\mu}n^{\nu}=0$) and $b$ is a small, dimensionless parameter measuring deviation from Lorentz invariance [^24] Notice that Bogoslovsky metric is an improper Lorentz norm, according to our definitions, as it is not smooth at the distinghished lightlike direction. Anyway, this means only that other specific geometric tools might be required for their treatment. Notice, for example, that if a power of $L$ is smooth outside the zero section as in [146], this is typically enough for the regularity of both, the cone and the indicatrix.. In 2007, VSR as well as its natural generalization to possibly curved spaces (called subsequently General Very Special Relativity), were clearly identified by Gibbons, Gomis and Pope as Finsler Geometry [70]; shortly after, Finslerian Cosmology was considered by Kouretsis et al. [106]. Fuster, Pfeifer and Pabst [67] considered Berwald spacetimes as a non-flat generalization of the line element used in VSR, and promoted the tidier name Very General Relativity. Pfeifer and Wohlfart [146] carried out a systematic construction of gravitational dynamics, including field equations (see §6.2.2), matter coupling, and observer definitions. Hohmann [85] explored relations between Finsler spacetimes and Cartan geometry.
@@ 6.1.2. The effective viewpoint
As pointed out above (§5.2.2, item (3)), anisotropic cones may emerge as an effective model for propagation coming from any fundamental theory. In this vein, Girelli, Liberati and Sindoni [73] noticed in 2007 that Planck-scale modifications to the usual energy-momentum relation (dispersion relations) in Quantum Gravity, can be understood by using Finsler geometry. These modifications are a common feature of all Quantum Gravity (QG) phenomenology approaches, and some models proposed that this phenomenology could be associated with an energy dependent geometry called “rainbow metric”. The authors show that rainbow metrics are naturally described by Finsler geometry, then providing a rigorous mathematical framework for the possible geometric structures arising in the semiclassical regime of QG. Moreover, Edwards and Kostelecký [49] connect effective field theories with Lorentz violation to Riemann-Finsler geometry and show that classical point-particle Lagrangians from Lorentz violating operators correspond to Finsler structures. Major cosmological issues as dark matter have been also modeled effectively by using Finsler elements, but the requirements for such a modelization may be disputed and go beyond our scope, see for example [41, 162].
@@ 6.1.3. The fundamental viewpoint
Following [22, Sect. 2-4] (which builds on [23, 89]), the theories for space and time such as Classical Mechanics and Special Relativity, rely on the existence of inertial frames of references (IFR). Such a frame $R$ assigns four coordinates $(t,x^{i})$ to the whole set of events. The first one $t$ is named temporal, the other three $x^{i}$ spatial, and they obey the following two postulates:
@@
By combining these two structures under certain compatibility conditions, EPS argues that the Lorentz metric of spacetime can be reconstructed —light rays and free-falling particles being model as lightlike and timelike pregeodesics— then providing the seeked foundation for General Relativity.
However, one can wonder why general Lorentz-Finsler metrics are excluded for EPS. Tavakol and Van den Berg [160] pointed out that a Lorentz norm (or, whith more generality a Berwald space) should satisfy EPS axioms, as their geodesics underlie an affine space. Lammërzhal and Perlick [110] argued against the unjustified requirement of smoothness for the radar coordinates introduced around each event $e$ in the EPS first axiom [^25] Specifically, the smoothness at $e$ of a function $g(p)=-t(e_{1})t(e_{2})$, meaning the product of local arrival parameters of two light rays (then satisfying $g(e)=0,g_{,a}(e)=0$)..
The careful analysis in [22, §5] identifies the two (unjustified) physical hypotheses on smoothness which turns out the exclusion of general Lorentz-Finsler metrics, essentially:
(i) The requirement of smoothness pointed out by Lammërzhal and Perlick would imply to ask for the smoothness at 0 of $F^{2}$, where $F$ is a Finsler metric, specifically, the one in a cone triple $(\Omega,T,F)$ associated with the cone $\mathcal{C}$ provided by the light rays see [22, Example 5.1]. By Lemma 2.2, this would make $F^{2}$ Riemannian and, then, $\mathcal{C}$ will be Lorentzian. Moreover, even in Lorentz-Minkowski spacetime there are situations where such a smoothness fail, see [22, footnote 24].
@@
(see [53, formula (7)]). Here, $\lambda$ depends on the parameterization $x^{a}(u)$ of the curve and $\Pi^{a}_{bc}$ ’s depend on $x^{a}$ and are called the projective coefficients (indeed, they determine a projective structure $\mathcal{P}$ compatible with some affine connection). However, $\Pi^{a}_{bc}$ ’s are not permitted to depend on the direction of the velocities $\dot{x}^{j}$, and this not justified, see [22, §5.2.2]. Indeed, if permitted would a non-linear connection would be obtained (its exponential as regular as explained in (5)).

### v2

@@ 5. Applications in Everyday Physics
    Finsler spacetime techniques can solve technical issues about propagation which are more involved from a purely Finslerian viewpoint (for example, see §5.2.1 below).
@@ 5.1. Fermat, Zermelo and Snell problems
Hamilton Jacobi equation and optical geometrics (see §2.5.4) are linked to the classical Fermat principle which states that the trajectories of light rays are the critical points of the arrival time functional, $\int n(\gamma(s))ds$, on curves $\gamma$ connecting the source with each point. This underlies the relativistic and classical applications below, where the time is embedded in the manifold structure even in the classical setting, providing a richer geometric picture.
@@ 5.1.1. An increasingly general Fermat principle
Figure 11. The first and last causal curves from $p$ to $\alpha$ (in green) must be lightlike geodesics by causality considerations —in higher dimensions other critical points of the arrival functional can be pictured.
@@
In the relativistic setting, the Fermat principle was posed by Kovner [116] and studied rigurously by Perlick [153]. This author also noticed that the existence of the global time $t$ was irrelevant and, indeed the parameter of $\alpha$ for any regular parametrization (i.e., smooth with non-vanishing velocity) could play its role for the functional $\mathcal{T}$. As the computations of critical points only uses variations of the arriving curve, global hypotheses in $\mathcal{C}$ are required only to ensure existence of critical points. Perlick also made a first extension of Fermat principle to the Lorentz-Finsler case [155].
Perlick and Piccione [156] also extended the principle to the case when the source of the lightlike curves is a spacelike submanifold $P$ (instead of a point $p$) and the curves arrive at a timelike submanifold (now actually endowed with a time function in order to redefine properly $\mathcal{T}$). In the setting of SSTK spacetimes above, Caponio et al. [41] considered also the case when the arrival curve $\alpha$ is not timelike, noticing that just the non-orthogonality between $\alpha$ and the incident curve $\gamma$ is sufficient to establish Fermat principle (notice that this condition is automatically satisfied when $\alpha$ is timelike), see [41, Th. 7.4].
@@ 5.1.2. Zermelo navigation
If $\mathcal{C}$ is a SSTK structure [^20] The following assertions could be extended to Finsler SSTK spacetimes, taking into account footnote 15. then its whole causality is encoded in a wind Finsler structure $\Sigma$ on a slice $S=\{t=0\}$ through $p$. Thus, the general conditions which ensure causal simplicity or global hyperbolicity for $\mathcal{C}$ can be expressed accurately in terms of $\Sigma$ as explained in §4.1. If $\alpha$ is an integral curve of the Killing vector field $K$ intersecting $S$ at a point $q$, then the existence of Zermelo solutions is equivalent to the existence of minimizing geodesics $\rho$ for the wind Finsler structure from $p$ to $q$. If, additionally, $\alpha$ is parametrized by the time $t$, then the arrival time is the wind Finsler length of $\rho$. In case that the SSTK is a standard stationary spacetime, the wind Finsler structure $\Sigma$ becomes just a Randers metric and $\rho$ is a usual geodesic. Moreover, as the integral curves of $K$ are timelike, the Fermat principle holds with no restriction about orthogonality at the arrival point in $\alpha$. From the geometric viewpoint, this generalizes widely the previous study by Perlick about the stationary case in [154], where the possibility of a reduction in one dimension less for Fermat principle considered.
@@ 5.1.3. Snell law
The basic Snell law for (anisotropic) cones is ([100, Th. 4.3], see Fig. 14):
@@
where $u$ (resp $v$) is the incident (resp. refracted) $\mathcal{C}_{1}$-lightlike (resp. $\mathcal{C}_{2}$-lightlike) direction at $o\in\eta$, $H_{u}^{1}$ (resp. $H_{v}^{2}$) is the hyperplane tangent to $\mathcal{C}_{1}$ at $u$ (resp. to $\mathcal{C}_{2}$ at $v$) and Snell’s formula states that the intersections $\Pi_{u}^{1},\Pi_{v}^{2}$ of these hyperplanes with the tangent space to the interface $T_{o}\eta$ must be equal. There is, however, a number of issues in relation to (25) and Fig. 14:
@@
  • The orientations in these hyperplanes (selected by the side where the corresponding cone lives) must be consistent, as in Fig. 14. This consistency is the additional hypothesis to ensure local minimization for Fermat principle and, then, to obtain a Snell geodesic i.e. a locally horismotic curve. Indeed, if the blue cone lived on the other side of $H_{v}^{2}$, then the broken lightlike geodesic could not minimize the arrival time (and could not be used to compute the causal future of $p$).
  • Classical effects as reflection and total reflection also appear as a consequence of the Fermat problem.
@@
    For example, when $\eta$ is spacelike for $Q_{1}$ and $Q_{2}$, the reflection is forbidden. Moreover, if the cones point out in consistently oriented directions at $\eta$, two refracted trajectories would appear, see Fig. 15. In this case, only one of them is minimizing (a fact determined again by orientations) providing then a Snell geodesic. In case of non-consistent orientation, no refracted one can appear, see Fig. 16.212121In these figures, past cones (which were not considered in $\mathcal{C}$) are depicted and filled for an easier comparison with usual relativistic diagrams.
@@
All these possibilities are analyzed in [100, §5], as well as the corresponding causality in such discontinuous media [100, §6]. About the latter, it is worth pointing out the higher technical difficulty of the Lorentz-Finsler case in comparison with the Lorentz one, stressed in the proof of [100, Lemma 6.6].
@@ 5.2.1. General wave propagation and wildfires
Reasoning as for Zermelo navigation in §5.1.2, let $(M,\mathcal{C})$ be a cone structure modelling the velocity of wave propagation in each event and direction permitted by a medium. In the case of Cosmology, the medium is vacuum, but in the case of a problem in classical Mechanics (the propagation of a wildfire or a sound wave), a cone triple for $(\Omega,T,F)$ on $\mathcal{C}$ appears naturally. Namely, $T$ would model the observers “at rest” on Earth, $\Omega$ is $dt$, with $t$ the Newtonian time, and $F$ (which depends on $t$) specifies $\mathcal{C}$.
Focusing on a wildfire, it will start at a bounded region $R$ which can be modelled as a compact open subset with boundary $S=\partial R$ of the (spacelike) hypersurface $t=0$. Then, the evolved burned region is equal to its $\mathcal{C}$-causal future $J^{+}(R)$. The spacetime frontfire will be its boundary in $t>0$, which matches the horismos $E^{+}(S)$. The frontfire at each instant $t_{0}>0$ is $E^{+}(S)\cap\{t=t_{0}\}$ (see Fig. 17). Now, Lorentz-Finsler geometry applies to show that $E^{+}(S)$ is composed by the cone geodesics starting at $S$ with initial velocity pointing outside $\mathds{R}$. Notice that, at each point $p\in S$, there are exactly two lightike directions orthogonal to $S$, the (non-depicted) second one pointing inwards $R$.
This general viewpoint of anisotropic wave propagation was developed in [101]. It includes the short time existence of solutions [101, Theorem 4.8]. This result is more involved now than in the purely Lorentz case (see [101, Lemma 4.7]), indeed, it had not been carried out previously (i.e., with a non-spacetime approach) as far as we know.
The particular case of wildfires is analyzed detailedly in [103]. This includes a comparison with actual wildfire monitorization, which is based in rough approximations to the wave equation. As emphasized therein [103, §5.2], focal and cut points along cone geodesics become points where wildfires are more intense and firefighters might get trapped.
@@ 5.2.2. From seisms to Numerics and Particle Physics
The modelization of seisms is similar to wildfires. However, it is natural to consider discontinuous layers for seisms and, thus, a systematic use of Snell formula. In this classical setting, the interface between different layers yields a timelike interface $\eta$ where the basic Snell formula (25) is enough. Indeed, typically, the interface will be “constant in time” so that its spacetime evolution is trivial [^22] Certainly, the effect of the seism itself could modify the interface, however, this would be negligible for the computation of the sismic wavefront. Moreover, in any case, it is not expected that the velocity of the movement of the interface could be compared with the velocity of seism propagation so that (25) would remain..
Anyway, the use of Snell formula can go far beyond the scope of seims, and require the use of other cases explained below (25). Specifically:
@@
    Computer aided modelizations of classical systems must use discretizations. In a natural way, this modelization yield a grid where each cell would contains a cone (Fig. 18). Then, the propagation of the wave between each two grids will be controlled by the basic Snell formula (25).
@@
    Such a discretization migth be useful for Numerical Relativity. The usual way to proceed in this field is considering at each cell the Lorentzian metric at one of its points. Notice, however, that such a Lorentzian metric is determined by the cone structure $\mathcal{C}$ up to a conformal factor. Thus, the separate study of $\mathcal{C}$ as above and this factor might be more efficient.
    For this purpose, the interface would be generically timelike or spacelike. The first case is essentially governed by Snell formula (25). For spacelike interfaces, on the one hand, the possibility of double refraction (as in Fig. 15) appears naturally. Then, the possibility to distinguish the Snell geodesics by using orientations would be useful. On the other hand, the case of no refraction (Fig. 16) would mean either that the grid approximation is overly rough or that a singularity appears.
@@ 6.1.1. Very Special and General Relativity
where $\eta_{\mu\nu}$ is the Minkowski metric, $n^{\mu}$ is a fixed lightlike direction ($\eta_{\mu\nu}n^{\mu}n^{\nu}=0$) and $b$ is a small, dimensionless parameter measuring deviation from Lorentz invariance [^24] Bogoslovsky metric is an improper Lorentz norm, according to our definitions, as it is not smooth at the distinghished lightlike direction. Anyway, this means only that other specific geometric tools might be required for their treatment. Notice also that if a power of $L$ is smooth outside the zero section (as in [159]), this is typically enough for the regularity of both, the cone and the indicatrix.. In 2007, VSR as well as its natural generalization to possibly curved spaces (called subsequently General Very Special Relativity), were clearly identified by Gibbons, Gomis and Pope as Finsler Geometry [74]; shortly after, Finslerian Cosmology was considered by Kouretsis et al. [114]. Fuster, Pfeifer and Pabst [71] studied Berwald spacetimes as a non-flat generalization of the line element used in VSR, and promoted the tidier name Very General Relativity. Pfeifer and Wohlfart [159] carried out a systematic construction of gravitational dynamics, including field equations (see §6.2.2), matter coupling, and observer definitions. Hohmann [92] explored relations between Finsler spacetimes and Cartan geometry.
@@ 6.1.2. The effective viewpoint
As pointed out above (§5.2.2, item (3)), anisotropic cones may emerge as an effective model for propagation coming from any fundamental theory. In this vein, Girelli, Liberati and Sindoni [77] noticed in 2007 that Planck-scale modifications to the usual energy-momentum relation (dispersion relations) in Quantum Gravity, can be understood by using Finsler geometry. These modifications are a common feature of many Quantum Gravity (QG) phenomenology approaches, and some models proposed that this phenomenology could be associated with an energy dependent geometry called “rainbow metric”. The authors showed that rainbow metrics are naturally described by Finsler geometry, then providing a mathematical framework for the possible geometric structures arising in the semiclassical regime of QG. Moreover, Edwards and Kostelecký [56] connected effective field theories with Lorentz violation to Finsler geometry and showed that classical point-particle Lagrangians from Lorentz violating operators correspond to Finsler structures. Major cosmological issues as dark matter have been also modeled effectively by using Finsler elements, but the requirements for such a modelization may be disputed and go beyond our scope, see for example [43, 177].
@@ 6.1.3. The fundamental viewpoint
Following [24, Sect. 2-4] (which builds on [25, 97]), the theories for space and time such as Classical Mechanics and Special Relativity, rely on the existence of inertial frames of references (IFR). Such a frame $R$ assigns four coordinates $(t,x^{i})$ to the whole set of events. The first one $t$ is named temporal, the other three $x^{i}$ spatial, and they obey the following two postulates:
@@
By combining these two structures under certain compatibility conditions, EPS argues that the Lorentz metric of spacetime can be reconstructed —light rays and free-falling particles being modelled as lightlike and timelike pregeodesics— then providing the seeked foundation for General Relativity.
However, one can wonder why general Lorentz-Finsler metrics are excluded for EPS. Tavakol and Van den Berg [175] pointed out that a Lorentz norm (or, whith more generality a Berwald space) should satisfy EPS axioms, as their geodesics underlie an affine space. Lämmerzahl and Perlick [118] argued against the unjustified requirement of smoothness for the radar coordinates introduced around each event $e$ in the EPS first axiom [^25] Specifically, the smoothness at $e$ of a function $g(p)=-t(e_{1})t(e_{2})$, meaning the product of local arrival parameters of two light rays (then satisfying $g(e)=0,g_{,a}(e)=0$)..
The careful analysis in [24, §5] identifies the two (unjustified) physical hypotheses on smoothness which turns out the exclusion of general Lorentz-Finsler metrics, essentially:
(i) The requirement of smoothness pointed out by Lämmerzahl and Perlick would imply to ask for the smoothness at 0 of $F^{2}$, where $F$ is a Finsler metric, specifically, the one in a cone triple $(\Omega,T,F)$ associated with the cone $\mathcal{C}$ provided by the light rays see [24, Example 5.1]. By Lemma 2.2, this would make $F^{2}$ Riemannian and, then, $\mathcal{C}$ will be Lorentzian. Moreover, even in Lorentz-Minkowski spacetime there are situations where such a smoothness fail, see [24, footnote 24].
@@
(see [57, formula (7)]). Here, $\lambda$ depends on the parameterization $x^{a}(u)$ of the curve and $\Pi^{a}_{bc}$ ’s depend on $x^{a}$ and are called the projective coefficients (indeed, they determine a projective structure $\mathcal{P}$ compatible with some affine connection). However, $\Pi^{a}_{bc}$ ’s are not permitted to depend on the direction of the velocities $\dot{x}^{j}$, and this not justified, see [24, §5.2.2]. Indeed, if permitted a non-linear connection would be obtained (its exponential being as regular as explained in (6)).